Tay-Sachs disease is an autosomal recessive disorder. Among Ashkenazi Jews of central European ancestry about in 1/3600 children are born with Tay-Sachs disease. What is the frequency for being a carrier for Tay-Sachs in this population? A carrier does not show the trait but can pass it on to his/her offspring.
Answer #1

92 = 1/3600 => 0.0002717 This , q => 0.0166 P = 1-0.0166 = 0.983 The frequency of carriers => 2(PXq). Answer => 0.0326hope it helps.